# Towards AI Search Paradigm

### v1

@@ 1 Introduction
In an age where individuals are constantly managing an overwhelming flow of data, Information Seeking (IS), an active behavior that constructs new cognition when faced with knowledge gaps (Wilson, 1981; Dervin et al., 1983), has emerged as a vital process for making informed decisions and solving complex problems. The advent of web search engines marked a significant leap in IS. Web search engines function as specialized systems built upon fundamental Information Retrieval (IR) technologies, systematically crawling, indexing, and retrieving internet-based information in response to user queries. Over the past decades, the field of IR has undergone transformative generational shifts to achieve superior performance. Researchers and practitioners have progressed systematically from early lexical models to machine learning paradigms, and ultimately, to large language models.
@@
Lexical IR technologies predominantly rely on keyword matching techniques, encompassing vector space models, probabilistic frameworks, and traditional language models (Brin and Page, 1998; Yang, 2005; Robertson et al., 2009; Mikolov, 2013). These approaches represent documents and queries as bags of words, estimating relevance based on the exact or partial overlap of terms. While effective for exact term matching, they inherently struggle with semantic mismatches and vocabulary variation—failing, for example, to associate synonymous expressions or contextual nuances. Such limitations have increasingly highlighted the need for more robust retrieval mechanisms. This motivated the field’s transition toward semantic-aware retrieval paradigms capable of deeper information assimilation.
@@
Learning-to-Rank (LTR) methodologies (Burges et al., 2005; Cao et al., 2007; Qin et al., 2010) emerged from the pursuit of more relevant and higher-quality search results, driving the application of machine learning to ranking problems. This next generation of information retrieval systems replaced heuristic ranking approaches with machine-learned models that directly optimize ranking objectives. LTR techniques leveraged feature engineering—incorporating signals including text-matching scores, document structure metadata, authority metrics, and user behavior signals (e.g., click-through rates)—combined with supervised machine learning to optimize relevance ranking. This represented a significant advancement in ranking effectiveness, consistently placing more relevant documents in top positions compared to lexical retrieval models. The output of such systems is a ranked list of documents, requiring users to perform an additional step to access the target information: clicking on relevant documents and then locating or synthesizing the needed content — revealing a gap between system outputs and true information needs.
@@
The advent of Large Language Models (LLMs) (Wei et al., 2022a; Achiam et al., 2023; Bubeck et al., 2023; Mesnard et al., 2024; Dubey et al., 2024) has enabled information retrieval systems to bridge this gap, transitioning from document retrieval to direct generation of precise contextualized answers. However, current RAG systems primarily function as single-shot answer generators and struggle with queries involving complex information needs, such as innovation-related inquiries, emotional nuance, or proactive planning, the orchestration of diverse tools, and deep reasoning across multiple, sometimes conflicting, pieces of evidence (Fan et al., 2024; Zhao et al., 2024). These systems’ output quality is critically dependent on the initial document retrieval phase, lacking both robustness in handling imperfect retrieval and the sophisticated reasoning capabilities needed to address truly multifaceted user needs. For instance, a seemingly straightforward question like "Who was older, Emperor Wu of Han or Julius Caesar, and by how many years?" poses significant challenges for current systems. While individual birth dates for both historical figures may be easily retrievable, no single document explicitly compares their ages. This requires the system to perform a complex, multi-stage process: (1) retrieving and verifying each emperor’s birth year from separate sources, (2) resolving potential conflicts between contradictory records, (3) calculating the age difference, and (4) synthesizing the final comparison. Such queries demand multi-step reasoning capabilities—the ability to decompose complex questions, execute sequential sub-queries, evaluate evidence reliability, and integrate intermediate results into a coherent answer.
@@
This paper establishes that the historical evolution of IR systems necessitates a fundamental advancement: cognitive architectures that genuinely emulate human information foraging behaviors and multi-stage reasoning processes. To meet this need, we propose a revolutionary IS paradigm, AI Search Paradigm — a collaborative, multi-agent framework powered by large language models that reasons, plans, and executes complex problem — solving strategies on the user’s behalf. Drawing inspiration from early work on computational algorithms (Knuth, 1997) and system interactions (Cohen et al., 2007), this AI search paradigm coordinates specialized agents, each responsible for a distinct phase of the information-seeking process, to jointly deliver accurate, context-rich answers that mirror human-level inquiry. The key agent roles include:
@@
  • Master Agent. Master serves as the initial point of contact, responsible for analyzing the user’s query to assess its complexity and intent. Based on the nature and difficulty of the problem, Master dynamically coordinates and assembles an appropriate team of subsequent agents. This is a unique feature of the proposed paradigm, compared to traditional IR systems, which incorporate static query understanding modules (Xiong et al., 2024), and RAG systems, where initial query analysis typically leads to a fixed processing pipeline (Ma et al., 2023). In addition, it continuously evaluates the performance of subordinate agents; in the event of task failures, Master conducts reflective analysis and guides the team to re-plan and re-execute accordingly.
@@
  • Planner Agent. Planner is only invoked for complex queries that require multi-step reasoning or information gathering. Given the user’s input query, it selects corresponding tools from a conceptualized Model-Context Protocol (MCP) Servers Platform (Edwards, 2025) to dynamically adjust LLMs’ capability boundary. Planner then decomposes the overarching query into a structured sequence of manageable sub-tasks, represented as a Directed Acyclic Graph (DAG). This step involves strategic planning of how to approach the problem. While advanced RAG systems might perform query decomposition or multi-hop reasoning (Ma et al., 2023), Planner’s function of generating a comprehensive, explicit DAG of sub-tasks with interdependencies and potential tool bindings represents a more flexible and proactive planning capability than typically found in standard RAG systems (Lewis et al., 2020b).
@@
  • Executor Agent. Executor is responsible for carrying out the simple query or the individual sub-tasks defined by Planner. A crucial aspect of its role is the invocation of external tools, selected by Planner from an MCP Servers Platform (Edwards, 2025), to gather information or perform computations necessary for each sub-task. It also evaluates the outcomes of these executions. Traditional IR systems execute internal algorithms, and while some RAG models can use tools, Executor’s role in managing a diverse set of tools for DAG-derived sub-tasks, including outcome assessment and fallback mechanisms, suggests a more robust and orchestrated execution layer (Qu et al., 2025b).
@@
  • Writer Agent. The final agent in the workflow, Writer, synthesizes the information and results gathered from all completed sub-tasks. Its goal is to generate a coherent, contextually rich, and potentially multi-perspective response to the user’s original query, including necessary filtering and disambiguation. While the generator component in RAG systems also synthesizes answers (Lewis et al., 2020b), Writer in the AI search paradigm is envisioned to work with a more structured and diverse set of inputs from various planned sub-tasks, aiming for a more comprehensive and conversational synthesis than generation from a flat list of retrieved documents.
@@
This paradigm emphasizes adaptability through dynamic team configurations and aims to achieve comprehensive and context-aware information synthesis, moving beyond simple document retrieval (Abril and Plant, 2007).
@@ 2 System Overview
In this section, we introduce a modular AI search paradigm designed to address the limitations of conventional retrieval-augmented systems in handling complex, multi-step information needs. Inspired by human collaborative search behavior, the framework coordinates multiple specialized agents to evaluate, plan, execute, and synthesize in a scalable and context-aware manner. Specifically, the agents are assigned a range of specialized responsibilities, including proactively evaluating the complexity of user queries, configuring appropriate multi-agent teams, conducting task decomposition and planning, selecting suitable tools via the MCP servers platform, executing the selected tools to fulfill specific sub-tasks, and synthesizing the outcomes of all sub-tasks to generate comprehensive and multi-perspective responses. Such a coordinated architecture ensures a methodical progression of the search process toward coherent and comprehensive outcomes.
@@
  • Planner. This agent is only invoked for complex queries that require multi-step reasoning or information gathering, and is responsible for task decomposition and strategic planning in response to complex queries, enabling the AI search system to address them through a sequence of simpler, structured sub-tasks. Its core function is to select corresponding tools from an MCP platform for dynamically adjusting LLMs’ capability boundary and generating a DAG that encodes a global task plan, where each node represents an atomic and schedulable sub-task, and each edge captures the conditional dependency between sub-tasks derived from the reasoning process of the Planner. Furthermore, when the Master evaluates sub-tasks as having execution errors or lacking critical data, the Planner, under the Master’s direction, undertakes a reconfiguration of the DAG.
@@
  • Executor-Inclusive Configuration. This configuration handles moderately complex queries that require factual information from external sources but do not involve multi-step reasoning or decomposition. The Master determines that the query can be resolved through single-step execution and assigns it directly to the Executor. Meanwhile, the AI search system uses the query semantics along with contextual cues derived from the team configuration to query the MCP servers platform, retrieving a focused and semantically relevant subset of tools. The Executor selects the top-ranked tool from this subset and invokes it to perform the task. Upon obtaining the result, it evaluates whether the output meets the task requirements. Once execution is complete, the output is forwarded to the Writer, who synthesizes the final response. For instance, given the query “Is Beijing’s weather good for going out today?”, the Executor invokes the retrieved weather query tool, obtains real-time weather data, and after verifying its completeness, passes the result to the Writer. The response might be: “Beijing’s weather today is sunny, with temperatures ranging from 12°C to 25°C. It is suitable for outdoor activities, although precautions against UV exposure are advised.
  • Planner-Enhanced Configuration. This configuration is designed for complex queries that require multi-step reasoning and structured task execution. When high query complexity is detected, the Master delegates control to the Planner, initiating a planning-based workflow. In parallel, tool selection by the Planner is supported by querying the MCP servers platform for a semantically filtered candidate set, based on the input query and current execution context. The Planner decomposes the query into a sequence of atomic sub-tasks and organizes them into a DAG that captures their logical and execution dependencies. For each sub-task, a suitable tool is selected from the retrieved tool subset and explicitly bound to the corresponding node. The Executor then traverses the DAG layer by layer, invoking the bound tools and evaluating the intermediate results. Once all sub-tasks are completed, the aggregated outputs are passed to the Writer, who synthesizes them into a coherent and context-aware final response. For example, consider a complex query “Who is elder, Emperor Han-Wu or Emperor Caesar, by how many years?”. In this scenario, the Master delegates the query to the Planner, who is responsible for both planning and decomposing it into three specific sub-tasks, each associated with a specific tool from the candidate set:
@@ 3.1 Background
When addressing complex queries that necessitate multi-step sub-task execution and the coordinated invocation of multiple tools, traditional RAG systems often fall short in effective task planning and dynamic multi-tool coordination. Consequently, their reliance on single-step task execution or fixed-tool utilization frequently results in incomplete information retrieval or inaccurate responses. We revisit the motivating example from a large-scale real-world search engine in Section 2 to underscore the limitations of traditional systems and the motivate the design of a Planner: “Who is elder, Emperor Han-Wu or Emperor Caesar, by how many years?”. To resolve this query, it is first decomposed into three specific sub-tasks, each associated with a dedicated tool. Then, these sub-tasks are organized into a DAG that reflects their logical execution order and subsequently executed in sequence. Throughout this section, we will employ this example to provide a comparative analysis of how recent RAG systems and our proposed search system approach the same problem.
@@
Moreover, advanced RAG approaches, as illustrated in the middle part of Fig. 2 (i.e., ReAct (Yao et al., 2023a)), introduce a structured loop that alternates between Thought, Action, and Observation phases. In such frameworks, the model plans the subsequent reasoning step, retrieves relevant information, and iteratively incorporates the results. When applied to the current case, a ReAct-like agent would first consider retrieving the birth dates, then carry out document retrievals for both emperors, and update its internal state accordingly. However, despite these enhancements, ReAct’s reliance solely on in-context memory precludes genuine external tool invocation; consequently, even if the agent correctly determines which emperor is older, it cannot precisely compute the age difference due to the absence of a dedicated calculator tool and structured numerical reasoning. Furthermore, the iterative context expansion inherent in ReAct complicates the maintenance of entity alignment and consistent task planning, thereby increasing the risk of error propagation. Similarly, RQ-RAG (Chan et al., 2024) seeks to enhance RAG performance by explicitly decomposing complex queries into sequential subqueries, solving them step-by-step while retaining intermediate results. For instance, a query decomposition-based agent would first retrieve Emperor Han-Wu’s birth date, then that of Emperor Caesar, and subsequently deduce their age difference. Nonetheless, despite the improved decomposition capabilities, RQ-RAG still operates within a single retrieval-and-generation framework that lacks dynamic tool utilization. As a result, although it may successfully obtain individual birth dates, it is unable to perform the subsequent computational reasoning necessary to determine the precise age difference, leaving the answer incomplete. In summary, while ReAct and RQ-RAG represent significant advancements over traditional RAG architectures by incorporating iterative reasoning and query decomposition, their inability to dynamically invoke external tools and to plan tasks restricts their effectiveness in accurately resolving complex, multi-step queries.
@@
Necessity of Task Planner. To overcome these limitations, a dedicated Planner is indispensable for complex query resolution. Unlike previous RAG systems that rely exclusively on retrieval or in-context decomposition, a Planner explicitly decomposes queries into fine-grained sub-tasks, determines their logical dependencies via a DAG, and dynamically selects appropriate tools for task execution beyond simple retrieval, as illustrated in the right part of Fig. 2. Moreover, the planner is capable of re-planning; if any intermediate results deviate from the expected objectives, the Planner, under the guidance of the Master, adjusts the task plan accordingly. These capabilities fundamentally expand the scope of retrieval-augmented systems, transforming passive “retrieve-then-generate” pipelines into “reason, plan, execute and re-plan”, a proactive AI search system. Concretely, the planner must decide what can be planned, which tools should be invoked, and how to optimize its own behaviour over time.

### v2

@@ 1 Introduction
In an age where individuals are constantly managing an overwhelming flow of data, Information Seeking (IS), an active behavior that constructs new cognition when faced with knowledge gaps [wilson1981user, dervin1983overview], has emerged as a vital process for making informed decisions and solving complex problems. The advent of web search engines marked a significant leap in IS. Web search engines function as specialized systems built upon fundamental Information Retrieval (IR) technologies, systematically crawling, indexing, and retrieving internet-based information in response to user queries. Over the past decades, the field of IR has undergone transformative generational shifts to achieve superior performance. Researchers and practitioners have progressed systematically from early lexical models to machine learning paradigms, and ultimately, to Large Language Models (LLMs).
@@
Lexical IR technologies predominantly rely on keyword matching techniques, encompassing vector space models, probabilistic frameworks, and traditional language models [brin1998anatomy, DBLP:journals/arist/Yang05, robertson2009probabilistic, mikolov2013efficient]. These approaches represent documents and queries as bags of words, estimating relevance based on the exact or partial overlap of terms. While effective for exact term matching, they inherently struggle with semantic mismatches and vocabulary variation—failing, for example, to associate synonymous expressions or contextual nuances. Such limitations have increasingly highlighted the need for more robust retrieval mechanisms. This motivated the field’s transition toward semantic-aware retrieval paradigms capable of deeper information assimilation.
@@
Learning-to-Rank (LTR) methodologies [RankNet:pairwise, ListNet:listwise, DBLP:journals/ir/QinLL10] emerged from the pursuit of more relevant and higher-quality search results, driving the application of machine learning to ranking problems. This next generation of information retrieval systems replaced heuristic ranking approaches with machine-learned models that directly optimize ranking objectives. LTR techniques leveraged feature engineering—incorporating signals including text-matching scores, document structure metadata, authority metrics, and user behavior signals (e.g., click-through rates)—combined with supervised machine learning to optimize relevance ranking. This represented a significant advancement in ranking effectiveness, consistently placing more relevant documents in top positions compared to lexical retrieval models. The output of such systems is a ranked list of documents, requiring users to perform an additional step to access the target information: clicking on relevant documents and then locating or synthesizing the needed content — revealing a gap between system outputs and true information needs.
@@
The advent of LLMs [wei2022chain, achiam2023gpt, bubeck2023sparks, DBLP:journals/corr/abs-2403-08295, DBLP:journals/corr/abs-2407-21783] has enabled information retrieval systems to bridge this gap, transitioning from document retrieval to direct generation of precise contextualized answers. However, current RAG systems primarily function as single-shot answer generators and struggle with queries involving complex information needs, such as innovation-related inquiries, emotional nuance, or proactive planning, the orchestration of diverse tools, and deep reasoning across multiple, sometimes conflicting, pieces of evidence [DBLP:conf/kdd/FanDNWLYCL24, DBLP:journals/corr/abs-2402-19473]. These systems’ output quality is critically dependent on the initial document retrieval phase, lacking both robustness in handling imperfect retrieval and the sophisticated reasoning capabilities needed to address truly multifaceted user needs. For instance, a seemingly straightforward question like "Who was older, Emperor Wu of Han or Julius Caesar, and by how many years?" poses significant challenges for current systems. While individual birth dates for both historical figures may be easily retrievable, no single document explicitly compares their ages. This requires the system to perform a complex, multi-stage process: (1) retrieving and verifying each emperor’s birth year from separate sources, (2) resolving potential conflicts between contradictory records, (3) calculating the age difference, and (4) synthesizing the final comparison. Such queries demand multi-step reasoning capabilities—the ability to decompose complex questions, execute sequential sub-queries, evaluate evidence reliability, and integrate intermediate results into a coherent answer.
@@
This paper establishes that the historical evolution of IR systems necessitates a fundamental advancement: cognitive architectures that genuinely emulate human information foraging behaviors and multi-stage reasoning processes. To meet this need, we propose a revolutionary IS paradigm, AI Search Paradigm — a collaborative, multi-agent framework powered by large language models that reasons, plans, and executes complex problem — solving strategies on the user’s behalf. Drawing inspiration from early work on computational algorithms [Knuth97] and system interactions [Cohen07], this AI search paradigm coordinates specialized agents, each responsible for a distinct phase of the information-seeking process, to jointly deliver accurate, context-rich answers that mirror human-level inquiry. The key agent roles include:
@@
  • Master Agent. Master serves as the initial point of contact, responsible for analyzing the user’s query to assess its complexity and intent. Based on the nature and difficulty of the problem, Master dynamically coordinates and assembles an appropriate team of subsequent agents. This is a unique feature of the proposed paradigm, compared to traditional IR systems, which incorporate static query understanding modules [xiong2024search], and RAG systems, where initial query analysis typically leads to a fixed processing pipeline [ma2023query]. In addition, it continuously evaluates the performance of subordinate agents; in the event of task failures, Master conducts reflective analysis and guides the team to re-plan and re-execute accordingly.
@@
  • Planner Agent. Planner is only invoked for complex queries that require multi-step reasoning or information gathering. Given the user’s input query, it selects appropriate tools from tool platforms—e.g., the conceptualized Model-Context Protocol (MCP) Servers Platform [edwards2025mcp]—to dynamically adjust the capability boundaries of the LLMs. Planner then decomposes the overarching query into a structured sequence of manageable sub-tasks, represented as a Directed Acyclic Graph (DAG). This step involves strategic planning of how to approach the problem. While advanced RAG systems might perform query decomposition or multi-hop reasoning [ma2023query], Planner’s function of generating a comprehensive, explicit DAG of sub-tasks with interdependencies and potential tool bindings represents a more flexible and proactive planning capability than typically found in standard RAG systems [DBLP:conf/nips/LewisPPPKGKLYR020].
@@
  • Executor Agent. Executor is responsible for carrying out the simple query or the individual sub-tasks defined by Planner. A crucial aspect of its role is the invocation of external tools, selected by Planner from tool platforms (e.g. MCP), to gather information or perform computations necessary for each sub-task. It also evaluates the outcomes of these executions. Traditional IR systems execute internal algorithms, and while some RAG models can use tools, Executor’s role in managing a diverse set of tools for DAG-derived sub-tasks, including outcome assessment and fallback mechanisms, suggests a more robust and orchestrated execution layer [quexploration].
@@
  • Writer Agent. The final agent in the workflow, Writer, synthesizes the information and results gathered from all completed sub-tasks. Its goal is to generate a coherent, contextually rich, and potentially multi-perspective response to the user’s original query, including necessary filtering and disambiguation. While the generator component in RAG systems also synthesizes answers [DBLP:conf/nips/LewisPPPKGKLYR020], Writer in the AI search paradigm is envisioned to work with a more structured and diverse set of inputs from various planned sub-tasks, aiming for a more comprehensive and conversational synthesis than generation from a flat list of retrieved documents.
@@
This paradigm emphasizes adaptability through dynamic team configurations and aims to achieve comprehensive and context-aware information synthesis, moving beyond simple document retrieval [Abril07].
@@ 2 System Overview
In this section, we introduce a modular AI search paradigm designed to address the limitations of conventional retrieval-augmented systems in handling complex, multi-step information needs. Inspired by human collaborative search behavior, the framework coordinates multiple specialized agents to evaluate, plan, execute, and synthesize in a scalable and context-aware manner. Specifically, the agents are assigned a range of specialized responsibilities, including proactively evaluating the complexity of user queries, configuring appropriate multi-agent teams, conducting task decomposition and planning, selecting suitable tools via tool platforms (e.g. the MCP servers platform), executing the selected tools to fulfill specific sub-tasks, and synthesizing the outcomes of all sub-tasks to generate comprehensive and multi-perspective responses. Such a coordinated architecture ensures a methodical progression of the search process toward coherent and comprehensive outcomes.
@@
  • Planner. This agent is only invoked for complex queries that require multi-step reasoning or information gathering, and is responsible for task decomposition and strategic planning in response to complex queries, enabling the AI search system to address them through a sequence of simpler, structured sub-tasks. Its core function is to select appropriate tools from a tool set (e.g. the MCP platform) to dynamically adjust the LLMs’ capability boundaries and to generate a DAG that encodes a global task plan. In this DAG, each node represents an atomic, schedulable sub-task, and each edge captures a conditional dependency between sub-tasks, derived from the Planner’s reasoning process. Furthermore, when the Master evaluates sub-tasks as having execution errors or lacking critical data, the Planner, under the Master’s direction, undertakes a reconfiguration of the DAG.
@@
  • Executor-Inclusive Configuration. This configuration handles moderately complex queries that require factual information from external sources but do not involve multi-step reasoning or decomposition. The Master determines that the query can be resolved through single-step execution and assigns it directly to the Executor. Meanwhile, the AI search system uses the query semantics along with contextual cues derived from the team configuration to query the tool platform server (e.g. MCP server), retrieving a focused and semantically relevant subset of tools. The Executor selects the top-ranked tool from this subset and invokes it to perform the task. Upon obtaining the result, it evaluates whether the output meets the task requirements. Once execution is complete, the output is forwarded to the Writer, who synthesizes the final response. For instance, given the query “Is Beijing’s weather good for going out today?”, the Executor invokes the retrieved weather query tool, obtains real-time weather data, and after verifying its completeness, passes the result to the Writer. The response might be: “Beijing’s weather today is sunny, with temperatures ranging from 12°C to 25°C. It is suitable for outdoor activities, although precautions against UV exposure are advised.
  • Planner-Enhanced Configuration. This configuration is designed for complex queries that require multi-step reasoning and structured task execution. When high query complexity is detected, the Master delegates control to the Planner, initiating a planning-based workflow. In parallel, tool selection by the Planner is supported by querying the tool platform server (e.g. MCP server) for a semantically filtered candidate set, based on the input query and current execution context. The Planner decomposes the query into a sequence of atomic sub-tasks and organizes them into a DAG that captures their logical and execution dependencies. For each sub-task, a suitable tool is selected from the retrieved tool subset and explicitly bound to the corresponding node. The Executor then traverses the DAG layer by layer, invoking the bound tools and evaluating the intermediate results. Once all sub-tasks are completed, the aggregated outputs are passed to the Writer, who synthesizes them into a coherent and context-aware final response. For example, consider a complex query “Who is elder, Emperor Han-Wu or Emperor Caesar, by how many years?”. In this scenario, the Master delegates the query to the Planner, who is responsible for both planning and decomposing it into three specific sub-tasks, each associated with a specific tool from the candidate set:
@@ 3.1 Background
When addressing complex queries that necessitate multi-step sub-task execution and the coordinated invocation of multiple tools, traditional RAG systems often fall short in effective task planning and dynamic multi-tool coordination. Consequently, their reliance on single-step task execution or fixed-tool utilization frequently results in incomplete information retrieval or inaccurate responses. We revisit the motivating example from a large-scale real-world search engine in Section 2 to underscore the limitations of traditional systems and motivate the design of the Planner: “Who is elder, Emperor Han-Wu or Emperor Caesar, by how many years?”. To resolve this query, it is first decomposed into three specific sub-tasks, each assigned to a dedicated tool. Then, these sub-tasks are organized into a DAG that reflects their logical execution order and subsequently executed in sequence. Throughout this section, we will employ this example to provide a comparative analysis of how recent RAG systems and our proposed search system approach the same problem.
@@
Moreover, advanced RAG approaches, as illustrated in the middle part of Fig. 2 (i.e., ReAct [yao2023react]), introduce a structured loop that alternates between Thought, Action, and Observation phases. In such frameworks, the model plans the subsequent reasoning step, retrieves relevant information, and iteratively incorporates the results. When applied to the current case, a ReAct-like agent would first consider retrieving the birth dates, then carry out document retrievals for both emperors, and update its internal state accordingly. However, despite these enhancements, ReAct’s reliance solely on in-context memory precludes genuine external tool invocation; consequently, even if the agent correctly determines which emperor is older, it cannot precisely compute the age difference due to the absence of a dedicated calculator tool and structured numerical reasoning. Furthermore, the iterative context expansion inherent in ReAct complicates the maintenance of entity alignment and consistent task planning, thereby increasing the risk of error propagation. Similarly, RQ-RAG [chan2024rq] seeks to enhance RAG performance by explicitly decomposing complex queries into sequential subqueries, solving them step-by-step while retaining intermediate results. For instance, a query decomposition-based agent would first retrieve Emperor Han-Wu’s birth date, then that of Emperor Caesar, and subsequently deduce their age difference. Nonetheless, despite the improved decomposition capabilities, RQ-RAG still operates within a single retrieval-and-generation framework that lacks dynamic tool utilization. As a result, although it may successfully obtain individual birth dates, it is unable to perform the subsequent computational reasoning necessary to determine the precise age difference, leaving the answer incomplete. In summary, while ReAct and RQ-RAG represent significant advancements over traditional RAG architectures by incorporating iterative reasoning and query decomposition, their inability to dynamically invoke external tools and to plan tasks restricts their effectiveness in accurately resolving complex, multi-step queries.
@@
Necessity of Task Planner. To overcome these limitations, a dedicated Planner is indispensable for complex query resolution. Unlike previous RAG systems that rely exclusively on retrieval or in-context decomposition, a Planner explicitly decomposes queries into fine-grained sub-tasks, determines their logical dependencies via a DAG, and dynamically selects appropriate tools for task execution beyond simple retrieval, as illustrated in the right part of Fig. 2. Moreover, the planner is capable of re-planning; if any intermediate results deviate from the expected objectives, the Planner, under the guidance of the Master, adjusts the task plan accordingly. These capabilities fundamentally expand the scope of retrieval-augmented systems, transforming passive “retrieve‐then-generate” pipelines into “reason, plan, execute and re-plan”, a proactive AI search system. Concretely, the planner must decide what can be planned, which tools should be invoked, and how to optimize its own behaviour over time.